# FlowDAS: A Flow-Based Framework for Data Assimilation

### v1

@@ 3.1 Stochastic Interpolants for Data Assimilation
3:     Compute $\tilde{\bm{I}}_{s}^{k}$ and $\bm{R}_{s}^{k}$ using (14) for $k\in\mathcal{B}_{K^{\prime}}$
4:     Compute the empirical loss $\mathcal{L}_{b}^{\text{emp}}(\hat{\bm{b}})$ in Equation 13
@@
Estimation of $\nabla\log p(\bm{y}\mid\bm{X}_{s},\bm{X}_{0})$
The term $\nabla\log p(\bm{y}\mid\bm{X}_{s},\bm{X}_{0})$ captures the observation information. Since the observation model only directly links $\bm{y}=\bm{y}_{k+1}$ and $\bm{X}_{1}=\bm{x}_{k+1}$, we compute this term by integrating with respect to $\bm{X}_{1}$:
$\nabla\log p(\bm{y}\mid\bm{X}_{s},\bm{X}_{0})$ $=\frac{\nabla\,p(\bm{y}|\bm{X}_{s},\bm{X}_{0})}{p(\bm{y}|\bm{X}_{s},\bm{X}_{0})}=\frac{\nabla\int p(\bm{y}|\bm{X}_{1})p(\bm{X}_{1}|\bm{X}_{s},\bm{X}_{0})\,\mathrm{d}\bm{X}_{1}}{\int p(\bm{y}|\bm{X}_{1})p(\bm{X}_{1}|\bm{X}_{s},\bm{X}_{0})\,\mathrm{d}\bm{X}_{1}}$ (10)
$\nabla\log p(\bm{y}\mid\bm{X}_{s},\bm{X}_{0})$ $\approx\frac{\nabla\,\frac{1}{J}\sum_{j=1}^{J}p(\bm{y}\mid\bm{X}_{1}^{(j)})}{\frac{1}{J}\sum_{j=1}^{J}p(\bm{y}\mid\bm{X}_{1}^{(j)})}=\frac{\sum_{j=1}^{J}p(\bm{y}\mid\bm{X}_{1}^{(j)})\,\nabla\,\log p(\bm{y}\mid\bm{X}_{1}^{(j)})}{\sum_{j=1}^{J}p(\bm{y}\mid\bm{X}_{1}^{(j)})}$ (11)
$=\sum_{j=1}^{J}w_{j}\nabla\,\log p(\bm{y}\mid\bm{X}_{1}^{(j)}),$
Acceleration of Monte Carlo sampling
Accurate sampling from $p(\bm{X}_{1}\mid\bm{X}_{s},\bm{X}_{0})$ requires solving the SDE in Equation 4, which can be computationally intensive. To improve efficiency, we propose using approximate numerical integration methods such as Euler’s method (Süli and Mayers, 2003, p. 317) or Heun’s method (Süli and Mayers, 2003, p. 324):
$\begin{split}\text{Euler: }\hat{\bm{X}}_{1}&=\bm{b}_{s}(\bm{X}_{s},\bm{X}_{0})(1-s)+\int\sigma_{s}\,\mathrm{d}\bm{W}_{s},\\ \text{Heun: }\hat{\bm{X}}_{1}^{\prime}&=\frac{\bm{b}_{s}(\bm{X}_{s},\bm{X}_{0})+\bm{b}_{1}(\hat{\bm{X}}_{1},\bm{X}_{0})}{2}(1-s)+\int\sigma_{s}\,\mathrm{d}\bm{W}_{s}.\end{split}$ (12)
These methods offer first-order and second-order approximations, respectively, introducing slight numerical bias but significantly accelerated sampling speed (i.e., $\mathcal{O}((1-s)^{2})$ for Euler and $\mathcal{O}((1-s)^{3})$ for Heun). In practice, Heun’s method consistently demonstrated better performance than Euler’s method. Section A.2 further compares these approximations.
@@
1:  Input: Observation $\bm{y}_{1:K}$, the measurement map $\mathcal{A}$, initial state $\bm{x}_{0}$, model $\hat{\bm{b}}_{s}(\bm{X},\bm{X}_{0})$, noise coefficient $\sigma_{s}$, grid $s_{0}=0<s_{1}<\cdots<s_{N}=1$, i.i.d. $\bm{z}_{n}\sim\mathcal{N}(0,\bm{I}_{D})$ for $n=0:N-1$, step size $\zeta_{n}$, Monte Carlo sampling times $J$
2:  Set $\hat{\bm{x}}_{0}\leftarrow\bm{x}_{0}$
@@
8:        $\{\hat{\bm{X}}^{(j)}_{1}\}^{J}_{j=1}\leftarrow$ Posterior estimation ($\hat{\bm{b}}_{s}$, $s_{n}$, $\bm{X}_{0}$, $\bm{X}_{s_{n}}$) by Equation 12
@@
14:  return  $\{\hat{\bm{x}}_{k}\}_{k=1}^{K}$
3.2 Implementation Details
Training
$\mathcal{L}_{b}^{\text{emp}}(\hat{\bm{b}})=\frac{1}{K^{\prime}}\sum_{k\in\mathcal{B}_{K^{\prime}}}\int_{0}^{1}\|\hat{\bm{b}}_{s}(\tilde{\bm{I}}_{s}^{k},\bm{x}_{k})-\bm{R}_{s}^{k}\|^{2}\,\,\mathrm{d}s,$ (13)
where $K^{\prime}\leq K$ and $\mathcal{B}_{K^{\prime}}\subset\{0:K\}$ is a subset of indices of cardinality $K^{\prime}$, with
$\tilde{\bm{I}}_{s}^{k}$ $=\alpha_{s}\bm{x}_{k}+\beta_{s}\bm{x}_{k+1}+\sqrt{s}\sigma_{s}\bm{z}_{k}$ (14)
[…]
Inference
Our inference procedure generates trajectories conditioned on observations $\bm{y}_{1:K}$ using the learned drift model $\hat{\bm{b}}_{s}(\bm{X}_{s},\bm{X}_{0})$. We start by setting a specific state $\bm{x}_{k}$ as initial $\bm{X}_{0}$ and iterate over a predefined temporal grid $s_{0}=0<s_{1}<\cdots<s_{N}=1$. Within each iteration, we first compute posterior estimates $\{\hat{\bm{X}}^{(j)}_{1}\}^{J}_{j=1}$ using Equation 12 for $J$ times. Then, we move one step further towards $s_{N}$ on $\bm{X}_{s_{n}}$ by solving Equation 9 which involves backpropagating the gradient $\nabla_{\bm{X}_{s_{n}}}\sum_{j=1}^{J}w_{j}\|\bm{y}-\mathcal{A}(\hat{\bm{X}}^{(j)}_{1})\|_{2}^{2}$ to enforce consistency with observations $\bm{y}_{k}$. We iterated this process autoregressively, by setting $\bm{X}_{s_{0}}^{k+1}=\bm{X}_{s_{N}}^{k}$. Empirically, we found that using a constant step size $\zeta_{n}$ across the inference process produced generally good results, although fine-tuning $\zeta_{n}$ at each step can slightly improve the performance (Bai et al., 2024; Chung et al., 2022). The chosen $J$ values, reported in Table S.6, were sufficient large for stable performance, with subtle variation across the dimensionality of problems. An ablation study, detailed in Section A.1, further confirmed that larger $J$ offers diminishing returns. Overall, Algorithm 2 summarizes the inference process.
4 Experiments and Results
This section evaluates our proposed framework, FlowDAS, on a range of tasks with a consistent parameter setting ($\alpha_{s}=1-s,\sigma_{s}=1-s,\beta_{s}=s^{2}$). We test FlowDAS on both low- and high-dimensional stochastic dynamical systems, including low-dimensional problems with high-order observation models, such as the double-well potential (Section B.2.5) and the chaotic Lorenz 1963 system, as well as high-dimensional tasks involving the incompressible Navier-Stokes equations. Additionally, we demonstrate its practical applicability in a realistic Particle Image Velocimetry (PIV) simulation. These results underscore the versatility and robustness of FlowDAS.
4.1 Lorenz 1963
In this experiment, we evaluate the performance of our FlowDAS framework using the Lorenz 1963 system, a widely studied benchmark in the DA community (Lorenz, 1963; Rozet and Louppe, 2023; Bao et al., 2023). The state vector of the Lorenz system, $\bm{x}=(a,b,c)\in\mathbb{R}^{3}$, evolves according to the following nonlinear stochastic ordinary differential equations (ODEs):
$\,\mathrm{d}a/\,\mathrm{d}t$ $=\mu(b-a)+\xi_{1},$ (15)
$\,\mathrm{d}b/\,\mathrm{d}t$ $=a(\rho-c)-b+\xi_{2},$
$\,\mathrm{d}c/\,\mathrm{d}t$ $=ab-\tau c+\xi_{3},$
where $\mu=10$, $\rho=28$, and $\tau=8/3$ define the ODE parameters, and $\bm{\xi}=(\xi_{1},\xi_{2},\xi_{3})\in\mathbb{R}^{3}$ is the process noise, with each component having a standard deviation $\sigma=0.25$. This chaotic system poses a significant challenge for numerical methods, so we use the fourth-order Runge-Kutta (RK4) method (Noar et al., 2024) to simulate its state transition (see Section B.3 for details).
$y=\mathcal{A}(\bm{x})+\bm{\eta}=\arctan(a)+\eta,$ (16)
Dataset
We generate 1,024 independent trajectories, each containing 1,024 states, and split the data into training (80%), validation (10%), and evaluation (10%) sets. Initial states are sampled from the statistically stationary regime of the Lorenz system, with additional data generation details provided in Section B.3. For this low-dimensional problem, we use a fully connected neural network to approximate the drift term in stochastic interpolants; the network architecture is also described in the same appendix.
Baselines and metrics
We compare our method against two baselines: the SDA solver with a fixed window size of 2 and the classic bootstrap particle filter (BPF) (Gordon et al., 1993). Section B.3 details the score network architecture for SDA and particle density settings for BPF.
We evaluate the performance of FlowDAS and baselines using four metrics: the expectation of log-prior $\mathbb{E}_{q(\bm{x}_{1:K}\mid\bm{y}_{1:K})}\left[\log p(\bm{x}_{2:K}\mid\bm{x}_{1})\right]$; the expectation of log data likelihood $\mathbb{E}_{q(\bm{x}_{1:K}\mid\bm{y}_{1:K})}\left[\log p(\bm{y}_{1:K}\mid\bm{x}_{1:K})\right]$; the Wasserstein distance (Villani, 2009) $W_{1}(\cdot,\cdot)$ between the true trajectory $\bm{x}_{1:K}$ and the estimated trajectory $\hat{\bm{x}}_{1:K}$; and the RMSE between the true and estimated states.
FlowDAS SDA BPF
$\uparrow$ $\log p(\hat{\bm{x}}_{2:K}\mid\hat{\bm{x}}_{1})$ 17.29 -332.7 17.88
$\uparrow$ $\log p(\bm{y}\mid\hat{\bm{x}}_{1:K})$0.228 -6.112 -1.572
$\downarrow$ $W_{1}(\bm{x}_{1:K},\hat{\bm{x}}_{1:K})$ 0.106 0.528 0.812
$\downarrow$ RMSE$(\bm{x}_{1:K},\hat{\bm{x}}_{1:K})$ 0.202 1.114 0.270
Table 1: Lorenz tracking results. This table summarizes the performance of FlowDAS, SDA, and BPF on the Lorenz 1963 experiment over 15 time steps. FlowDAS outperforms SDA across all evaluation metrics and is competitive with BPF, despite BPF utilizing the true transition equations, which are unknown to FlowDAS and SDA. The best results for each metric are highlighted in bold.
Results
We independently estimate 64 trajectories over 15 time steps using FlowDAS and the baseline methods. As shown in Table 1 and Figure 2, FlowDAS outperforms SDA across all metrics. FlowDAS is only slightly less effective than BPF in the expected log-prior, as BPF directly incorporates the true system dynamics into its state estimation.
The success of FlowDAS is primarily due to its accurate mapping from current to future states ($\bm{x}_{k}\rightarrow\bm{x}_{k+1}$). Despite lacking explicit transition equations, FlowDAS effectively captures the system dynamics through stochastic interpolants, enabling a closer approximation of state trajectories compared to SDA, which models joint distributions across sequential states using diffusion models. Additionally, stochastic interpolants allow FlowDAS to produce accurate state estimates while managing inherent variability, avoiding over-concentration on high-probability regions, and effectively dealing with rare events. This advantage is further illustrated in the double well potential experiment (Section B.2.5) where FlowDAS outperforms BPF, because BPF tends to be trapped by high-probability point estimates.
4.2 Incompressible Navier-Stokes Flow
Figure 3: Data assimilation of incompressible Navier-Stokes flow. The positive values (red) of the state, i.e., vorticity field, indicate clockwise rotation and negative values (blue) indicate counter-clockwise rotation. FlowDAS achieved results with more details and higher accuracy than the SDA both in the super-resolution task and the sparse observation task, showing FlowDAS’s efficiency in tackling DA tasks with highly non-linear complex systems. Additionally, FlowDAS is also better at recovering high-frequency information, evidenced by the spectral analysis in Figure S.9.
where $\bm{\omega}$ represents the vorticity field, the state variable in this fluid dynamics system ($\bm{x}=\bm{\omega}$). The velocity $\bm{v}=\nabla^{\perp}\psi=(-\partial_{y}\psi,\partial_{x}\psi)$ is expressed in terms of the stream function $\psi(x,y)$, which satisfies $-\Delta\psi=\bm{\omega}$. The term $d\bm{\xi}$ represents white-in-time random forcing acting on a few Fourier modes, with parameters $\nu,\alpha,\varepsilon>0$ specified in Section B.4.1.
$y=\mathcal{A}(\bm{\omega})+\bm{\eta},$ (18)
Dataset
In this experiment, system dynamics are simulated by solving Equation 17 using a pseudo-spectral method (Peyret, 2002) with a resolution of $256^{2}$ and a timestep $\Delta t=10^{-4}$. We simulate 200 flow conditions over $t\in[0,100]$, saving snapshots of fluid vorticity field ($\bm{\omega}=\nabla\times v$) at the second half of each trajectory ($t\in[50,100]$) at intervals of $\Delta t=0.5$ with a reduced resolution of $128^{2}$. The data are divided into training (80%), validation (10%), and evaluation (10%) sets.
Baselines and metrics
We compare our method against a SDA solver with a fixed window size of 2. BPF is not included in our testing, as its particle requirements grow exponentially with system dimensions, making it impractical for high-dimensional fluid dynamics systems. Additional details on the model architecture and training for both SDA and FlowDAS are provided in Section B.4.3.
We evaluate performance using the RMSE between the predicted and ground-truth vorticity fields. Additionally, we assess the reconstruction of the kinetic energy spectrum to determine whether the physical characteristics of the fluid are accurately preserved. Section B.4.5 and Figure S.9 provide the definitions and results for the kinetic energy spectrum metric.
$32^{2}\rightarrow 128^{2}$ $16^{2}\rightarrow 128^{2}$ $5\%$ $1.5625\%$
FlowDAS 0.038 0.067 0.071 0.123
SDA 0.073 0.133 0.251 0.258
[…]
Results
We conduct experiments across different observation resolutions ($32^{2}$, $16^{2}$) and observation sparsity levels (5%, 1.5625%). For the super-resolution task, the goal is to reconstruct high-resolution vorticity data ($128^{2}$) from low-resolution observations. In the inpainting task, only 5% or 1.5625% of pixel values are retained, with the rest set to zero, and we attempt to recover the complete vorticity field. The model is evaluated on four unseen datasets (2$\times$ super-resolution, 2$\times$ inpainting), with 64 samples for each configuration. Quantitative metrics, including RMSE and kinetic energy spectrum comparisons, are provided in Table 2 and Figure S.9.
Figure 3 presents a visual comparison between FlowDAS and SDA for the flow super-resolution and inpainting tasks. Our method consistently outperforms SDA in terms of reconstruction accuracy and resolution, capturing high-frequency information with greater precision. This advantage is further validated by the kinetic energy spectrum in Figure S.9. The RMSE scores in Table 2 further highlight the effectiveness of FlowDAS in accurately estimating the underlying fluid dynamics from observational data.
4.3 Particle Image Velocimetry
This section presents a realistic application of our method: Particle Image Velocimetry (PIV). PIV is a widely used optical technique for measuring velocity fields in fluids, with many scientific applications in aerodynamics (Taylor et al., 2010; Van Oudheusden, 2013; Koschatzky et al., 2011), biological flow studies (King et al., 2007; Stamhuis, 2006; Ergin et al., 2018), and medical research (Chen et al., 2014; Tan et al., 2009; Özcan et al., 2023).
In a standard PIV setup, as shown in Figure 4, fluorescent tracer particles are seeded into a fluid flowing through a channel with transparent walls. A laser sheet illuminates the fluid, and particle movements are recorded by a high-speed camera with adjustable temporal resolution. By analyzing the displacement of these tracer particles, the velocity field within the fluid can be determined at sparse locations. Unlike the task in Section 4.2, which involves recovering dense vorticity fields from sparse vorticity observations, PIV introduces a slightly different DA task: recovering dense vorticity fields ($\bm{x}=\bm{\omega}$) from sparse velocity measurements. This observation model is defined by
$y=\mathcal{A}(\bm{v}(\bm{\omega}))+\bm{\eta},$ (19)
Figure 4: Illustration of the real-world Particle Image Velocimetry (PIV) experiment: The flow is seeded with tracer particles illuminated by a laser sheet, and their movements are captured by a camera to derive the sparse velocity field. The goal here is to recover dense vorticity field from the sparse velocity measurement.
Dataset
In this experiment, we used the same fluid dynamics simulation data from the NS experiments described in Section 4.2. However, we convert the vorticity data to velocity fields via Fourier transform to create synthetic PIV datasets (Lagemann et al., 2021). The particle positions in our simulation were randomly initialized and then perturbed according to the simulated flow motion pattern. In these synthetic images, we assumed a particle density of 0.03 particles per pixel, a particle diameter of 3 pixels, and a peak intensity of 255 for each particle in grayscale. The images were processed through a standard PIV pipeline to extract particle locations, match corresponding particles across frames, and compute sparse velocity observations. These sparse measurements were then used in DA to reconstruct the full vorticity fields.
Baselines and metrics
As in Section 4.2, we compared our method against the SDA solver with a fixed window size of 2. Instead of training new scores or stochastic interpolant networks, we directly adopted the networks trained on vorticity data from the incompressible NS flow experiment to evaluate FlowDAS and SDA on the PIV task. The same quantitative metrics from Section 4.2 were employed to validate the performance of each method.
Figure 5: Data assimilation of Particle Image Velocimetry. The vorticity field is visualized in the same way as in Figure 3. FlowDAS significantly outperforms SDA in terms of reconstruction resolution and RMSE, recovering more detailed features even when direct observations are not available. These improvements highlight the potential of FlowDAS for real-world applications.
Results
FlowDAS accurately recovers the underlying fluid motion from the observed particle images, yielding vorticity estimates with minimal error relative to the ground truth. As shown in Figure 5 and Section B.5.3, FlowDAS significantly outperformed SDA in terms of reconstruction resolution and RMSE. This improvement underscores FlowDAS’s robustness and applicability in fluid dynamics research.
5 Conclusion and Future Work
In this work, we introduced FlowDAS, a flow-based data assimilation framework designed to address the challenges of high-dimensional, nonlinear dynamical systems. By leveraging stochastic interpolants, FlowDAS effectively integrates complex transition dynamics with observational data, enabling accurate state estimation without relying on explicit physical simulations. Through experiments on both low- and high-dimensional systems—including the Lorenz 1963 system, incompressible Navier-Stokes flow, and Particle Image Velocimetry—FlowDAS demonstrated strong performance in recovering accurate state variables from sparse, noisy observations. These results highlight FlowDAS as a robust alternative to traditional model-driven methods (e.g., particle filters) and data-driven approaches (e.g., score-based data assimilation), offering improved accuracy, efficiency, and adaptability. Future work will focus on scaling the framework to more complex real-world systems (Huang et al., 2024; Seabra et al., 2024) and further optimizing its performance with Sim2Real setting (Chen et al., 2024a).
Appendix A Ablation Study
@@
A.1 Monte Carlo Sampling and An Alternative
In Equation 10 and Equation 11, we estimate $\nabla\log p\,(\bm{y}\mid\bm{X}_{s},\bm{X}_{0})$ by
@@ A.1 Monte Carlo Sampling and An Alternative
A.2 The Method For Posterior Estimation
We also evaluate the impact of different methods for posterior estimation: Euler’s method (1st-order estimation) and Heun’s method (2nd-order estimation), as defined in Equation 12. The results are presented in Table S.5, where “1st-order" and “2nd-order" refer to Euler’s and Heun’s methods, respectively. "No correction" indicates forecasting purely based on the model without incorporating observation information. The results show that both 1st-order and 2nd-order estimations provide reasonable accuracy. However, the 2nd-order estimation (Heun’s method) consistently delivers better performance. This suggests that employing more accurate estimations of $p\,(\bm{X}_{1}|\bm{X}_{s},\bm{X}_{0})$ can effectively enhance model performance. Beyond the 2nd-order method, higher-order approaches like the Runge-Kutta 4th-order (RK4) method could further improve accuracy. However, these methods come with increased computational cost: 2nd-order estimation requires two neural network evaluations per step compared to one for 1st-order estimation, while RK4 requires four evaluations per step. In our experiments, we find that the 2nd-order estimation strikes a good balance between performance and efficiency, making it a practical choice. Further exploration of higher-order methods will be left for future research.
1st-order 2rd-order No correction

### v2

@@ B.1 Algorithms
3:     Compute $\tilde{\bm{I}}_{s}^{k}$ and $\bm{R}_{s}^{k}$ using (S.25) for $k\in\mathcal{B}_{K^{\prime}}$
4:     Compute the empirical loss $\mathcal{L}_{b}^{\text{emp}}(\hat{\bm{b}})$ in Equation S.24
@@
1:  Input: Observation $\bm{y}_{L:K}$, the measurement map $\mathcal{A}$, initial state $\bm{x}_{0}$, model $\hat{\bm{b}}_{s}(\bm{X},\bm{X}_{0})$, noise coefficient $\sigma_{s}$, grid $s_{0}=0<s_{1}<\cdots<s_{N}=1$, i.i.d. $\bm{z}_{n}\sim\mathcal{N}(0,\bm{I}_{D})$ for $n=0:N-1$, step size $\zeta_{n}$, Monte Carlo sampling times $J$
2:  Set $\hat{\bm{x}}_{L-1}\leftarrow\bm{x}_{L-1}$
@@
8:        $\{\hat{\bm{X}}^{(j)}_{1}\}^{J}_{j=1}\leftarrow$ Posterior estimation ($\hat{\bm{b}}_{s}$, $s_{n}$, $\bm{X}_{0}$, $\bm{X}_{s_{n}}$)
@@
14:  return  $\{\hat{\bm{x}}_{k}\}_{k=L}^{K}$
B.2 Constructing Training Dataset
Figure S.5: Structure of training data. Consecutive states are paired across multiple simulated trajectories to construct the $\tilde{\bm{I}}_{s}$ and $\bm{R}_{s}$ defined in Equation S.24 for training the velocity model $\bm{b}_{s}$.
  1. 1.
  2. 2.
    1. (a)
    2. (b)
  3. 3.
  4. 4.
    1. (a)
      sample states from $\bm{x}_{0:K-1}^{1:T}$ as $\bm{X}_{0}$’s.
    2. (b)
      retrieve the counterpart states from $\bm{x}_{1:K}^{1:T}$ as $\bm{X}_{1}$’s.
    The batches of state pairs $(\bm{X}_{0},\bm{X}_{1})$ are used to construct $\tilde{\bm{I}}_{s}$ and $\bm{R}_{s}$ in Equation S.24 for training the velocity model $\bm{b}_{s}$ as described in Algorithm 1.
In summary, we draw many state pairs $(\bm{X}_{0},\bm{X}_{1})$ and estimate the integral over $s$ in Equation S.24 by approximated via an empirical expectation over draws of $s\sim U(0,1)$, and for every $s$ and every state pairs $(\bm{X}_{0},\bm{X}_{1})$, we independently compute Equation S.25 with $S$ samples of $\bm{z}_{k}$, so the training loss can be rewritten as:
Monte Carlo Sampling Times $J$ Sampling Step Size $\zeta$
Double-well 17 1
Lorenz 1963 21 0.0002
Incompressible Navier Stokes SR 4x 25 1
SR 8x 25 2
SO 5% 25 1
SO 1.5625% 25 1.75
Particle Image Velocimetry 25 1
Weather Forecasting 25 0.1
B.3 FlowDAS Inference Hyperparameters
We present key hyperparameters for FlowDAS in the inference stage, including the Monte Carlo sampling times $J$ and the sampling step size $\xi$ in Table S.4.
Appendix C Ablation Study
@@
C.1 Monte Carlo Sampling and An Alternative
In Section A.3, we estimate $\nabla\log p\,(\bm{y}\mid\bm{X}_{s},\bm{X}_{0})$ by
@@ C.1 Monte Carlo Sampling and An Alternative
C.2 Posterior Estimation Methods
We also evaluate the impact of different methods for posterior estimation: as defined in Equations 9 and 10. The results are presented in Table S.7, where “1st-order” and “2nd-order” refer to 1st-order Milstein method and 2nd-order stochastic Runge-Kutta method, respectively. “No correction” indicates forecasting purely based on the model without incorporating observation information. The results show that both 1st-order and 2nd-order estimations provide reasonable accuracy. However, the 2nd-order estimation consistently delivers better performance. This suggests that employing more accurate estimations of $p\,(\bm{X}_{1}|\bm{X}_{s},\bm{X}_{0})$ can effectively enhance model performance. Beyond the 2nd-order method, higher-order approaches like the Runge-Kutta 4th-order (RK4) method could further improve accuracy. However, these methods come with increased computational cost: 2nd-order estimation requires two neural network evaluations per step compared to one for 1st-order estimation, while RK4 requires four evaluations per step. In our experiments, we find that the 2nd-order estimation strikes a good balance between performance and efficiency, making it a practical choice. Further exploration of higher-order methods will be left for future research.
1st-order 2nd-order No correction